OpenStudy (aaronandyson):

quadrilateral has the angle ratio 2 :3:3:4 find the angles in degrees and radians

OpenStudy (aaronandyson):

@Qwertty123

OpenStudy (aaronandyson):

I got 60,90,90 and 120 degrees.

rebeccaxhawaii (rebeccaxhawaii):

360° is the sum, your missing 40º.

OpenStudy (aaronandyson):

60+90 = 150 90+120 = 210 150+210 = 360

OpenStudy (aaronandyson):

whats missing?

rebeccaxhawaii (rebeccaxhawaii):

oui sorry. calculated wrong.